OpenStudy (cindy):

simplify 4/13 - 4/13 is the correct answer 8/13

OpenStudy (anonymous):

4/13 - 4/13=0

OpenStudy (anonymous):

bottom numbers are the same, so simply subtract the top numbers to get 0/13 which equals zero

OpenStudy (anonymous):

However, 4/13 - (-4/13)=8/13